# Frank–Wolfe algorithm for DC optimization problem

### v2

@@ Abstract
We propose two versions of the Frank–Wolfe (conditional gradient) algorithm with adaptive step sizes for DC optimization problems, where the objective is the difference of a differentiable convex function with Lipschitz continuous gradient and a convex (possibly non-differentiable) function. The second version builds on the first, using finite differences to approximate the gradient of the first component. Unlike previous approaches relying on curvature or Lipschitz constants, our stepsizes require no such parameters. We also introduce the class of piecewise-star-convex functions, which may be nonconvex and non–differentiable, and show that the first version achieves an $1/k$ convergence rate for minimizing such functions, while ensuring stationarity of all limit points. In the second version, the finite–difference scheme yields relative rather than absolute gradient errors, simplifying the analysis; we prove that all limit points are stationary and that the duality gap decreases at rate $1/\sqrt{k}$.
Keywords: Frank-Wolfe method; DC optimization problem; finite difference; weak-star-convex function.
@@ 1 Introduction
The DC optimization method involves minimizing a function that can be represented as the difference between two convex functions. We are interested in finding a solution to a constrained DC optimization problem, where the constraint set ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a convex and compact set, and f:=ghf:=g-h where $g:\mathbb{R}^{n}\to\mathbb{R}$ is a continuously differentiable convex function, and $h:\mathbb{R}^{n}\to\mathbb{R}$ is a convex function possibly non-differentiable. To the best of our knowledge, DC optimization can be traced back to pioneering works such as [42, 43], which introduced the first algorithms for this problem. Since then, the DC optimization has attracted the attention of the mathematical programming community (see for example [7, 11, 13, 32, 44]), not only for its own sake but also because it is an abstract model for several families of practical optimization problems. Although we are not concerned with practical issues at this time, we emphasize that practical applications emerge whenever the natural structure of the problem is modelled as a DC optimization problem, such as sparse generalized eigenvalue problems [41], sparse optimization problems [19], facility location and clustering problems [36].
@@
The Frank-Wolfe algorithm has a long history, dating back to Frank and Wolfe’s work in the 1950s to minimize convex quadratic functions over compact polyhedral sets, see [15]. This method was generalized about ten years later to minimize convex differentiable functions with Lipschitz continuous gradients and compact constraint convex sets, see [33]. Since then, this method has attracted the attention of several researchers who work with continuous optimization, thus becoming also known as the conditional gradient method. One of the factors that explain the interest in this method is its simplicity and ease of implementation. In fact, each method iteration only requires access to a linear minimization oracle over a compact convex set. It is also worth noting that, due to the method’s simplicity, it allows for low storage costs and ready exploration of separability and sparsity, making its application in large-scale problems quite attractive. It is interesting to note that the popularity of this method has increased significantly in recent years as a result of the emergence of several applications in machine learning, see [25, 29, 30]. For all of these reasons, several variants of this method have arisen throughout the years and new properties of it have been discovered, resulting in a large literature on it, papers dealing with this method include [4, 8, 9, 17, 18, 21, 27, 31, 34].
@@
In the present paper, we formulate two versions of Frank–Wolfe algorithm or conditional gradient method to solve DC optimization problem with an adaptive step size. The second version is based on the first and employs finite differences to approximate the gradient $\nabla g$. It is worth mentioning that the Frank-Wolfe algorithm has previously been formulated in the context of DC programming, see [26]. See also [46] which establishes theoretical connections between DC algorithms and convex-concave procedures with the Frank-Wolfe algorithm. In contrast to the previous study, which assume that the curvature/Lipschitz-type constant of f=ghf=g-h is bounded from above, the analysis of the Frank-Wolfe method done here just assumes that the gradient of the first component $g$ of $f$ is Lipschitz continuous. While designing the methods, even if we assume that the gradient of $g$ is Lipschitz continuous in both formulations, we will not compute the step size using the Lipschitz constant of $\nabla g$, as in earlier formulations that employed the curvature/Lipschitz-type constant of the objective function $f$. The step size will be computed adaptively, based on an idea introduced in [3] (see also [5, 39]) that approximates the Lipschitz constant. It has been shown in previous works that Frank-Wolfe algorithm produces a stationary point with a convergence rate of only ${\cal O}(1/\sqrt{k})$ for non-convex objective functions, see [28] (see also [26]). We introduce the concept of weak-star-convex function notion, which generalizes the star-convex function concept proposed in [37] as well as a few other related concepts, such as [22, 45]. The weak-star-convex objective functions are also taken into account in the convergence analysis of the first method presented here. The convergence rate is $\mathcal{O}(1/k)$ for both the function values and the duality gap, even though piecewise-star-convexs are generally non-differentiable and non-convex. Consequently, the class of functions for which the Frank–Wolfe method achieves an $\mathcal{O}(1/k)$ convergence rate includes piecewise-star-convexs that are differences of convex functions, with the first component having a Lipschitz continuous gradient. This class complements the set of star–convex functions, which was recently shown in [35] to enjoy the same convergence rate. As mentioned previously, the second version of the Frank-Wolfe algorithm uses finite difference to approximate the gradient of the function $g$. Similar to the strategies adopted in [20], the finite difference utilized here to approximate the gradient is computed with the step-size updated adaptively utilizing two previous iterations. Finite difference is an old concept that has been used to approximate derivatives in optimization settings (for example, see [38, Section 8.1]) and has already appeared in the study of the Frank-Wolfe algorithm (for example, see [14, 16, 25, 40]). It is worth noting that all previous applications of finite difference in the Frank-Wolfe algorithm provided approximate gradients with absolute error, whereas the one used here has the advantage of providing us with a relative error and thus being quite simple to analyze. We show that all limit points of the generated sequence for the second version of the Frank-Wolfe algorithm are stationary points for the problem under consideration, and we establish that the rate of convergence is ${\cal O}(1/\sqrt{k})$ for the duality gap.
@@
The remainder of the paper is organized as follows. Section 2 introduces the notation and collects auxiliary results. In Section 3 we define the concept of piecewise–star–convex, present its main properties, and provide illustrative examples. Section 4 states the optimization problem, the required assumptions, and related notation. Section 5 is devoted to the first version of the Frank–Wolfe algorithm, including its formulation, well–posedness, and analysis of the generated sequence. Section 6 presents the second version of the algorithm together with its formulation, well–posedness, and convergence analysis. Finally, Section 7 summarizes our conclusions.
@@ 2 Preliminaries
In this section, we recall some notations, definitions and basics results used throughout the paper. A function $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ is said to be convex if $\varphi(\lambda x+(1-\lambda)y)\leq\lambda\varphi(x)+(1-\lambda)\varphi(y)$, for all $x,y\in{\mathbb{R}^{n}}$ and $\lambda\in[0,1]$. And $\varphi$ is strictly convex when the last inequality is strict for $x\neq y$, for a comprehensive study of convex function see [23]. We say that $f:\mathbb{R}^{n}\to\mathbb{R}$ is locally Lipschitz if, for all $x\in\mathbb{R}^{n}$, there exist a constant ${\cal C}_{x}>0$ and a neighborhood $U_{x}$ of $x$ such that $|f(x)-f(y)|\leq{\cal C}_{x}\|x-y\|$, for all $y\in U_{x}.$ It is well known that, if $f:\mathbb{R}^{n}\to\mathbb{R}$ is convex, then $f$ is locally Lipschitz. And we also know that if $f:\mathbb{R}^{n}\to\mathbb{R}$ is continuously differentiable then $f$ is locally Lipschitz, see [10, p. 32]. Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a locally Lipschitz function. The Clarke’s subdifferential of $f$ at $x\in\mathbb{R}^{n}$ is given by $\partial_{c}f(x)=\{v\in\mathbb{R}^{n}:~f^{\circ}(x;d)\geq v^{{\scriptscriptstyle\mathrm{T}}}d,~\forall d\in\mathbb{R}^{n}\},$ where $f^{\circ}(x;d)$ is the generalized directional derivative of $f$ at $x$ in the direction $d$ given by
@@
For an extensive study of locally Lipschitz functions and Clarke’s subdifferential see [10, p. 27]. If $f$ is convex, then $\partial_{c}f(x)$ coincides with the subdifferential $\partial f(x)$ in the sense of convex analysis, and $f^{\circ}(x;d)$ coincides with the usual directional derivative $f^{\prime}(x;d)$; see [10, p. 36]. We recall that if $f:\mathbb{R}^{n}\to\mathbb{R}$ is differentiable, then $\partial_{c}f(x)=\{\nabla f(x)\}$ for any $x\in\mathbb{R}^{n}$; see [10, p. 33].
Theorem 2.1 ([10, p. 27]).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a locally Lipschitz function. Then, $\partial_{c}f(x)$ is a nonempty, convex, compact subset of $\mathbb{R}^{n}$ and $\|v\|\leq{\cal C}_{x},$ for all $v\in\partial_{c}f(x)$, where ${\cal C}_{x}>0$ is the Lipschitz constant of $f$ around $x$. Moreover, $f^{\circ}(x;d)=\max\{v^{{\scriptscriptstyle\mathrm{T}}}d:~v\in\partial_{c}f(x)\}$.
Theorem 2.2 ([10, p. 38-39]).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be given by f=ghf=g-h, where $g,h:\mathbb{R}^{n}\to\mathbb{R}$ is locally Lipschitz functions and $g$ is differentiable. Then, $f^{\circ}(x;d)=\nabla g(x)^{{\scriptscriptstyle\mathrm{T}}}d-h^{\prime}(x;d)$, for all $x,d\in\mathbb{R}^{n}$ and $\partial_{c}f(x)=\{\nabla g(x)\}-\partial h(x)$.
The next result is a combination of Theorem 2.2 with [10, Corollary on p. 52].
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a locally Lipschitz function and ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a closed and convex set. If $x^{*}\in{{\cal C}}$ is a minimizer of $f$ in $\cal C$}}, then $v^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, for all $v\in\partial_{c}f({x^{*}})$ and all $x\in{{\cal C}}$. As a consequence, if f=ghf=g-h with $g,h:\mathbb{R}^{n}\to\mathbb{R}$ is locally Lipschitz functions and $g$ is differentiable, then $(\nabla g({x^{*}})-u)^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, for all $u\in\partial_{c}h({x^{*}})$ and all $x\in{{\cal C}}$.
Hence, according to Theorem 2.3, every point satisfying the following inequality $v^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, for all $v\in\partial_{c}f({x^{*}})$ and for all $x\in{\cal C}$, is called to be a stationary point of $\min_{x\in{{\cal C}}}f(x)$. A continuously differentiable function $g:\mathbb{R}^{n}\to\mathbb{R}$ has gradient $\nabla g$ is $L$-Lipschitz continuous on ${{\cal C}}\subset{\mathbb{R}}^{n}$ if there exists a Lipschitz constant $L>0$ such that $\|\nabla g(x)-\nabla g(y)\|\leq L\|x-y\|$ for all   $x,y\in{{\cal C}}$. Thus, by using the fundamental theorem of calculus, we obtain the following result whose proof can be found in [6, Proposition A.24], see also[12, Lemma 2.4.2].
@@
Proposition 2.6 ([23, Proposition 6.2.1]).
Let $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be convex. Let $(x^{k})_{k\in{\mathbb{N}}}$ and $(u^{k})_{k\in{\mathbb{N}}}$ be sequences such that $u^{k}\in\partial h(x^{k})$, for all $k\in\mathbb{N}$. If $\lim_{k\to+\infty}x^{k}={\bar{x}}$ and $\lim_{k\to+\infty}u^{k}={\bar{u}}$, then ${\bar{u}}\in\partial h({\bar{x}})$.
Proposition 2.7 ([23, Proposition 6.2.2]).
@@
In the following we recall an useful results for our study on iteration complexity bounds for Frank–Wolfe algorithm, its proof can be found in [2, Lemma 13.13, Ch. 13, p. 387].
@@
3 Piecewise-star-convex functions
In this section, we introduce the concept of piecewise-star-convexity, a relaxation of the classical star–convexity notion from [37]. Many DC functions that are nonconvex nevertheless satisfy this weaker property. As will be shown in a later section, by identifying and exploiting piecewise–star–convexity we can broaden the class of problems for which the classical Frank–Wolfe algorithm attains an ${\cal O}(1/k)$ convergence rate. We begin by recalling the definition of star–convexity from [37] and then present its relaxed counterpart, the piecewise-star-convex notion, which will be central to our analysis.

### v3

# Frank–Wolfe algorithms for piecewise star-convex functions with a nonsmooth difference-of-convex structure

@@ Abstract
We study Frank–Wolfe algorithms for minimizing a nonconvex function expressed as the difference of a differentiable convex function with a Lipschitz continuous gradient and a possibly nonsmooth convex function over a compact convex set. We propose an adaptive Frank–Wolfe scheme that exploits this difference-of-convex structure and prove a sufficient descent property together with the convergence of the Frank–Wolfe gap to zero. We also introduce the notion of piecewise star-convexity, and show that for this class of functions Clarke-stationary points are minimizers on the cells of the underlying partition and, for finite partitions, we establish $\mathcal{O}(1/k)$ complexity bounds in function value and in the Frank–Wolfe gap. In particular, these bounds match the standard rates obtained in the convex case, showing that Frank–Wolfe methods can retain optimal complexity guarantees beyond the classical convex setting, even in the presence of nonsmooth terms.
Keywords: Frank-Wolfe algorithm; DC optimization problem; finite difference; piecewise-star-convex function.
@@ 1 Introduction
The DC optimization method involves minimizing a function that can be represented as the difference between two convex functions. We are interested in finding a solution to a constrained DC optimization problem, where the constraint set ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a convex and compact set, and f:=ghf:=g-h where $g:\mathbb{R}^{n}\to\mathbb{R}$ are continuously differentiable convex functions, and $h:\mathbb{R}^{n}\to\mathbb{R}$ is a convex function possibly non-differentiable. To the best of our knowledge, DC optimization can be traced back to pioneering works such as [37, 38], which introduced the first algorithms for this problem. Since then, the DC optimization has attracted the attention of the mathematical programming community (see for example [8, 12, 14, 29, 39]), not only for its own sake but also because it is an abstract model for several families of practical optimization problems. Although we are not concerned with practical issues at this time, we emphasize that practical applications emerge whenever the natural structure of the problem is modelled as a DC optimization problem, such as sparse generalized eigenvalue problems [36], sparse optimization problems [18], facility location and clustering problems [33].
@@
The Frank-Wolfe algorithm has a long history, dating back to Frank and Wolfe’s work in the 1950s to minimize convex quadratic functions over compact polyhedral sets, see [15]. This method was generalized about ten years later to minimize convex differentiable functions with Lipschitz continuous gradients and compact constraint convex sets, see [30]. Since then, this method has attracted the attention of several researchers who work with continuous optimization, thus becoming also known as the conditional gradient method. One of the factors that explains the interest in this method is its simplicity and ease of implementation. In fact, each method iteration only requires access to a linear minimization oracle over a compact convex set. It is also worth noting that, due to the method’s simplicity, it allows for low storage costs and ready exploration of separability and sparsity, making its application in large-scale problems quite attractive. It is interesting to note that the popularity of this method has increased significantly in recent years as a result of the emergence of several applications in machine learning, see [22, 25, 28]. For all of these reasons, several variants of this method have arisen throughout the years and new properties of it have been discovered, resulting in a large literature on it, papers dealing with this method include [4, 9, 10, 16, 17, 19, 24, 27, 31].
@@
In the present paper, we formulate a version of Frank–Wolfe algorithm or the conditional gradient method to solve DC optimization problem with an adaptive step size. It is worth mentioning that the Frank-Wolfe algorithm has previously been formulated in the context of DC programming, see [23]. See also [41] which establishes theoretical connections between DC algorithms and convex-concave procedures with the Frank-Wolfe algorithm. In contrast to the previous study, which assume that the curvature/Lipschitz-type constant of f=ghf=g-h is bounded from above, the analysis of the Frank-Wolfe method done here just assumes that the gradient of the first component $g$ of $f$ is Lipschitz continuous. While designing the methods, even if we assume that the gradient of $g$ is Lipschitz continuous, we will not compute the step size using the Lipschitz constant of $\nabla g$. The step size will be computed adaptively, based on an idea introduced in [3] (see also [5, 35]) that approximates the Lipschitz constant. It has been shown in previous works that Frank-Wolfe algorithm produces a stationary point with a convergence rate of only ${\cal O}(1/\sqrt{k})$ for non-convex objective functions, see [26] (see also [23]). Here, we introduce the concept of piecewise star-convex function notion, which generalizes the star-convex function concept proposed in [34] as well as a few other related concepts, such as [20, 40]. The convergence analysis developed in this paper covers piecewise star–convex with the nonsmooth difference-of-convex structure aforementined, we show that the proposed Frank–Wolfe scheme equipped with our backtracking stepsize attains the same order of iteration complexity as in the convex case, both the function-value and the Frank-Wolfe gap decay at the standard sublinear rate $\mathcal{O}(1/k)$. This holds even though piecewise star–convex functions are, in general, nonconvex and may be nonsmooth across cell boundaries. Consequently, our theory extends convex-type complexity guarantees for Frank–Wolfe methods to a broader nonconvex setting, including piecewise star–convex DC objectives with a smooth convex term $g$ and a convex, possibly nonsmooth term $h$. These guarantees complement recent results for star–convex functions in [32], and indicate that projection-free Frank–Wolfe schemes can maintain the usual first-order complexity behavior beyond classical convexity, even in the presence of nonsmooth terms.
@@
The remainder of the paper is organized as follows. Section 2 introduces the notation and collects auxiliary results. Section 3 presents the optimization problem, the standing assumptions, and related notation. In Section 4 we define piecewise star–convexity, establish its main properties, and provide illustrative examples. Section 5 is devoted to the proposed Frank–Wolfe algorithm, including its formulation, well-posedness, and the basic properties of the generated sequence. Section 5.1 analyzes the convergence properties of the method, and Section 5.2 derives the corresponding iteration–complexity bounds. Finally, Section 6 summarizes our conclusions.
@@ 2 Preliminaries
In this section, we recall some notations, definitions and basic results used throughout the paper. A function $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ is said to be convex if $\varphi(\lambda x+(1-\lambda)y)\leq\lambda\varphi(x)+(1-\lambda)\varphi(y)$, for all $x,y\in{\mathbb{R}^{n}}$ and $\lambda\in[0,1]$. And $\varphi$ is strictly convex when the last inequality is strict for $x\neq y$, for a comprehensive study of convex functions see [21]. We say that $f:\mathbb{R}^{n}\to\mathbb{R}$ is locally Lipschitz if, for all $x\in\mathbb{R}^{n}$, there exist a constant ${\cal C}_{x}>0$ and a neighborhood $U_{x}$ of $x$ such that $|f(x)-f(y)|\leq{\cal C}_{x}\|x-y\|$, for all $y\in U_{x}.$ It is well known that, if $f:\mathbb{R}^{n}\to\mathbb{R}$ is convex, then $f$ is locally Lipschitz. And we also know that if $f:\mathbb{R}^{n}\to\mathbb{R}$ is continuously differentiable then $f$ is locally Lipschitz, see [11, p. 32]. Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a locally Lipschitz function. The Clarke’s subdifferential of $f$ at $x\in\mathbb{R}^{n}$ is given by $\partial^{c}f(x)=\{v\in\mathbb{R}^{n}:~f^{\circ}(x;d)\geq v^{{\scriptscriptstyle\mathrm{T}}}d,~\forall d\in\mathbb{R}^{n}\},$ where $f^{\circ}(x;d)$ is the generalized directional derivative of $f$ at $x$ in the direction $d$ given by
@@
For an extensive study of locally Lipschitz functions and Clarke’s subdifferential see [11, p. 27]. If $f$ is convex, then $\partial^{c}f(x)$ coincides with the subdifferential $\partial f(x)$ in the sense of convex analysis, and $f^{\circ}(x;d)$ coincides with the usual directional derivative $f^{\prime}(x;d)$; see [11, p. 36]. We recall that if $f:\mathbb{R}^{n}\to\mathbb{R}$ is differentiable, then $\partial^{c}f(x)=\{\nabla f(x)\}$ for any $x\in\mathbb{R}^{n}$; see [11, p. 33].
Theorem 2.1 ([11, p. 27]).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a locally Lipschitz function. Then, $\partial^{c}f(x)$ is a nonempty, convex, compact subset of $\mathbb{R}^{n}$ and $\|v\|\leq{\cal C}_{x},$ for all $v\in\partial^{c}f(x)$, where ${\cal C}_{x}>0$ is the Lipschitz constant of $f$ around $x$. Moreover, $f^{\circ}(x;d)=\max\{v^{{\scriptscriptstyle\mathrm{T}}}d:~v\in\partial^{c}f(x)\}$.
Theorem 2.2 ([11, p. 38-39]).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be given by f=ghf=g-h, where $g,h:\mathbb{R}^{n}\to\mathbb{R}$ is locally Lipschitz functions and $g$ is differentiable. Then, $f^{\circ}(x;d)=\nabla g(x)^{{\scriptscriptstyle\mathrm{T}}}d-h^{\prime}(x;d)$, for all $x,d\in\mathbb{R}^{n}$ and $\partial^{c}f(x)=\{\nabla g(x)\}-\partial h(x)$.
The next result is a combination of Theorem 2.2 with [11, Corollary on p. 52].
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a locally Lipschitz function and ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a closed and convex set. If $x^{*}\in{{\cal C}}$ is a minimizer of $f$ in $\cal C$}}, then there exists $v\in\partial^{c}f(x^{*})$ such that $v^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, for all $x\in{{\cal C}}$. As a consequence, if f=ghf=g-h with $g,h:\mathbb{R}^{n}\to\mathbb{R}$ is locally Lipschitz functions and $g$ is differentiable, then there exists $u\in\partial^{c}h({x^{*}})$ such that $(\nabla g({x^{*}})-u)^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, all $x\in{{\cal C}}$.
Hence, according to Theorem 2.3, every point satisfying the following inequality $v^{{\scriptscriptstyle\mathrm{T}}}(x-{x^{*}})\geq 0$, for some $v\in\partial^{c}f({x^{*}})$ and for all $x\in{\cal C}$, is called to be a stationary point of $\min_{x\in{{\cal C}}}f(x)$. A continuously differentiable function $g:\mathbb{R}^{n}\to\mathbb{R}$ has gradient $\nabla g$ is $L$-Lipschitz continuous on ${{\cal C}}\subset{\mathbb{R}}^{n}$ if there exists a Lipschitz constant $L>0$ such that $\|\nabla g(x)-\nabla g(y)\|\leq L\|x-y\|$ for all   $x,y\in{{\cal C}}$. Thus, by using the fundamental theorem of calculus, we obtain the following result whose proof can be found in [7, Proposition A.24], see also[13, Lemma 2.4.2].
@@
Proposition 2.6 ([21, Proposition 6.2.1]).
Let $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be convex. Let $\{x^{k}\}_{k\in\mathbb{N}}$ and $(u^{k})_{k\in{\mathbb{N}}}$ be sequences such that $u^{k}\in\partial h(x^{k})$, for all $k\in\mathbb{N}$. If $\lim_{k\to+\infty}x^{k}={\bar{x}}$ and $\lim_{k\to+\infty}u^{k}={\bar{u}}$, then ${\bar{u}}\in\partial h({\bar{x}})$.
Proposition 2.7 ([21, Proposition 6.2.2]).
@@
In the following we recall an useful results for our study on iteration complexity bounds for Frank–Wolfe algorithm, its proof can be found in [6, Lemma 13.13, Ch. 13, p. 387].
@@
3 The DC optimization problem
$\begin{array}[]{c}\min_{x\in{{\cal C}}}f(x):=g(x)-h(x),\end{array}$ (1)
  • (A)
    $\|\nabla g(x)-\nabla g(y)\|\leq L\|x-y\|$ for all   $x,y\in{{\cal C}}$.
Since $\cal C$} is compact and $f$ is continuous on $\cal C$}, the problem (1) admits a finite global optimal value
$f_{\cal C}^{*}:=\min_{x\in{\cal C}}f(x)\;>\;-\infty,$ (3)
and the set of minimizers ${\cal C}^{*}:=\arg\min_{x\in{\cal C}}f(x)$ is nonempty. By Theorem 2.3, a point $\bar{x}\in{\cal C}$ is said to be Clarke stationary for problem (1) if there exists $\bar{u}\in\partial^{c}h(\bar{x})$ such that
$\bigl(\nabla g(\bar{x})-\bar{u}\bigr)^{\top}(x-\bar{x})\;\geq\;0,\qquad\forall\,x\in{\cal C}.$ (4)
In general, (4) is a necessary condition for optimality but may fail to be sufficient. In particular, every minimizer $x^{*}\in{\cal C}^{*}$ satisfies (4).
Let us introduce the linear minimization oracle and the associated gap function that will be used to define the Frank–Wolfe step for the DC model (1). Since f=ghf=g-h may be nonsmooth, the linearization is built from $\nabla g(x)$ and a subgradient of $h$. Given $x\in{\cal C}$ and $u\in\partial h(x)$, define
$p(x,u)\in\arg\min_{v\in{\cal C}}\bigl(\nabla g(x)-u\bigr)^{\top}(v-x),\qquad\omega(x,u):=\bigl(\nabla g(x)-u\bigr)^{\top}\bigl(p(x,u)-x\bigr).$ (5)
When no ambiguity arises, we write $p(x)$ and $\omega(x)$ for $p(x,u)$ and $\omega(x,u)$, where $u$ is the algorithmic choice at $x$.
Proposition 3.1.
Let ${\cal C}\subset\mathbb{R}^{n}$ be compact and convex. For every $x\in{\cal C}$ and every $u\in\partial^{c}h(x)$, let $p(x,u)$ and $\omega(x,u)$ be defined by (5). Then $\omega(x,u)\leq 0.$ Moreover, if $\omega(x,u)=0$, then $x$ is a stationary point for (1). Conversely, if $x\in{\cal C}$ is stationary, then there exists $u\in\partial^{c}h(x)$ such that $\omega(x,u)=0$.
Proof.
Fix $x\in{\cal C}$ and $u\in\partial^{c}h(x)$. Since the point $v=x$ is feasible in (5), we have
$\omega(x,u)=\min_{v\in{\cal C}}\bigl(\nabla g(x)-u\bigr)^{\top}(v-x)\leq\bigl(\nabla g(x)-u\bigr)^{\top}(x-x)=0,$
which proves $\omega(x,u)\leq 0$.
Furthermore, by (4), there exists $\bar{u}\in\partial^{c}h(\bar{x})$ such that $w(x,u)\geq 0$ if and only if $x$ is Clarke stationary, which completes the proof. ∎
We conclude this section with a standard example of a nonsmooth DC function, see a particular instance in [1, Example 2.4.]. It satisfies our standing assumptions. In particular, it admits a decomposition f=ghf=g-h in which $g$ is convex, differentiable, and has a Lipschitz continuous gradient, while $h$ is convex and possibly nonsmooth.
Example 3.2.
Let $A\in\mathbb{R}^{n\times n}$ have full column rank, so that $Q:=A^{\top}A$ is positive definite. Let $b\in\mathbb{R}^{n}$ and $\beta>0$. Define $f:{\mathbb{R}}^{n}\to\mathbb{R}$ by
$f(x)=\tfrac{1}{2}\,\|Ax-b\|_{2}^{2}-\beta\|x\|_{1}.$
Then $f$ is a nonsmooth DC function with the decomposition f=ghf=g-h, where $g(x):=\tfrac{1}{2}\,\|Ax-b\|_{2}^{2}$ and $h(x):=\beta\|x\|_{1}$. Moreover, $g$ is convex and differentiable with a Lipschitz continuous gradient, while $h$ is convex and nonsmooth.
Beyond its DC structure, this example illustrates a geometric feature that will be central in the sequel. Although $f$ is nonconvex, it becomes well behaved when restricted to suitable convex regions. Indeed, the sign pattern of $x$ induces a natural partition of $\mathbb{R}^{n}$ into orthants, and on each orthant the function reduces to a strictly convex quadratic. Consequently, any Clarke-stationary point is a minimizer of $f$ on its corresponding cell, even if it is not a global minimizer on $\mathbb{R}^{n}$. This motivates the class of piecewise star-convex models introduced in the next section, where star-type inequalities hold on each cell of a convex partition and can be used to obtain sharper convergence and complexity guarantees for Frank–Wolfe methods.
4 Piecewise star-convexity with a nonsmooth DC structure
In this section, we introduce the notion of piecewise star-convexity, a relaxation of the classical concept of star convexity from [34]. Although a difference-of-convex function is generally nonconvex, it often exhibits this weaker geometric property when restricted to suitable regions of the domain. As we show later, identifying and exploiting piecewise star-convexity yields sublinear convergence guarantees for Frank–Wolfe methods. Moreover, when the domain admits a finite partition, we obtain an ${\cal O}(1/k)$ rate on each region that matches the standard rate of the convex case, despite moving beyond the classical convex setting and allowing nonsmooth terms.
We begin by recalling the definition of star convexity from [34] and then introduce its piecewise counterpart, which will play a central role in our analysis.